1- If tax revenue is $400 billion and outlays are $600 billion, then
   a- there is a budget surplus of $200 billion
    b- there is a budget surplus of $1000 billion
    c- there is a budget deficit of $200 billion
    d- there is a budget deficit of $1000 billion

2- If equilibrium RGDP is $10000 billion and full employment GDP is $9000 billion, and the MPC = 0.75, what is the appropriate fiscal policy to return the economy to full employment GDP
    a- reduce taxes by $250 billion
    b- reduce taxes by $333.33 billion
    c- raise taxes by $250 billion
    d- raise taxes by $333.33 billion

3- The Prime Minister of Richlandia hires you as an economic consultant. He is concerned that the output level in Richlandia is too low and that this will cause unemployment to rise. He feels that it is necessary to increase output by $20 billion. He tells you that the MPC = 0.75. Which of the following would be the best advice to give to the Richlandia Prime Minister?
    a- increase government purchases by $15 billion
    b- reduce taxes by $10 billion
    c- reduce taxes by $4 billion
    d- increase government purchases by $5 billion


4- Which of the following statements is correct? I. The deficit and debt amounts must be presented as a percentage of GDP in order to have any meaning. II. The government borrows money through issuing shares of stock.
   a- I only
    b- II only
    c- both I and II
    d- neither I nor II

5- A minimal guideline for national debt is
   a- the debt should grow no faster than nominal GDP
    b- the debt should grow no faster than the budget deficit
    c- the debt should grow no faster than the decrease in the tax rate
    d- the debt should grow faster than nominal GDP
Answer #1

1. Option C

Explanation: There is budget deficit is the tax revenue is lower than the outlays.

2. Option D

Explanation: There is an inflationary gap of $10,000 - $9,000 = $1,000. The tax multiplier = -MPC/(1-MPC) = -0.75/(1-0.75) = -0.75/0.25 = -3